Part 1) A=1/2bh solve for H
Part 2)If the area of the triangle is 48cm(squared) and its base measurs 12cm, what is the height of the triangle?

Answer by nyc_function(2741)

Part (a)

A = 1/2*bH

We can write the right side as bH/2.

A = bH/2

2A = bH


2A/b = H

=====================================

Part (2)

A = bH/2

48*2 = 12H/2

96 = 12H

96/12 = H

8 = H

The height is 8cm.